OpenStudy (anonymous):

If n^-2 = 1/16, then n could be which of the following? -8 -4 -1/4 1/4 4

OpenStudy (anonymous):

\[n^{-2}=\frac{1}{n^2}=\frac{1}{16}\] go from there